# Integrable modules over quantum symmetric pair coideal subalgebras

### v1

@@ 1.1. Representation theory
It has been pointed out that the representation theory of $\imath$quantum groups develops very slowly ([Le19], [W21]). In fact, we do not still know any general theory to classify the finite-dimensional simple modules over $\imath$quantum groups of finite type, although such $\imath$quantum groups are quantum deformations of the universal enveloping algebra of a finite-dimensional complex reductive Lie algebra. Since the $\imath$quantum groups have close resemblance to the quantum groups, it seems to be essential for better understanding of their representation theory to formulate the notion of integrable modules over $\imath$quantum groups. This is the aim of the present paper.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 24K16903.
@@ 3.1. Quantum symmetric pair
Let $(I,I_{\bullet},\tau)$ be a generalized Satake diagram ([RV20, Definition 1]). From now on, we assume that the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$ is a Satake datum of type $(I,I_{\bullet},\tau)$ ([W23c, Definition 3.1.3]). In particular, the $I_{\bullet}$ is a subdatum of $I$ which is supposed to be of finite type, the $\tau$ is an involutive automorphism on the Cartan datum $I$, and the lattices $Y$ and $X$ are supposed to be equipped with involutive automorphisms $\tau$ which are compatible with the one on $I$. Let $w_{\bullet}\in W$ denote the longest element of the Weyl group $W_{\bullet}\subseteq W$ associated with $I_{\bullet}$. Set
@@ 3.2. Weight modules
A $\mathbf{U}^{\imath}$-module $V$ is said to be a weight module ([W23a, Definition 3.3.2]) if it admits a linear space decomposition
@@
as $\mathbf{U}^{\imath}$-modules. Then, the assertion can be found in [W23a, after Definition 3.3.5]. ∎
@@ 3.3. Integrable modules
Combining the two homomorphisms above, we see that $\mathbf{L}v^{\imath}_{\lambda,\mu}\simeq\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. By Lemma 3.3.2, the $\mathbf{U}$-module structure of $V^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}v^{\imath}_{\lambda,\mu}$. On the other hand, by [W23b, Proposition 3.4.3], the $\mathbf{U}$-module structure of $L^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. Hence, the assertion follows. ∎
@@ 4.1.1. Case $1:$ $\tau(k)=k=w_{\bullet}k$
Now, let us consider the $(n+1)$-dimensional irreducible representation $\mathbf{V}_{n}$ of $U_{q}(\mathfrak{sl}_{2})$, which we shall identify with the subalgebra of $\mathbf{U}$ generated by $E_{k}$, $F_{k}$, and $K_{k}^{\pm 1}$. By a similar argument to [W21, §3.1], one can show that the $B_{k}\in\mathbf{U}^{\imath}$ acts on $\mathbf{V}_{n}$ semisimply with eigenvalues $\kappa^{(n)},\kappa^{(n-2)},\dots,\kappa^{(-n)}$, where we set $\kappa^{(0)}:=\kappa_{k}$ and $b:=q_{k}\varsigma_{k}$. Therefore, we see that
@@ 4.2. A Presentation of the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$
$\{\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu}+\mathrm{wt}^{\imath}(b)}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}bm^{\imath}_{\lambda,\mu}\mid k\in I_{\circ},\ b\in\mathbf{B}(-\infty)\}.$
4.3. Integrable $\mathbf{U}$-modules
@@ 5. Quantized coordinate algebras
For example, the last two conditions above are satisfied if the parameters ${\boldsymbol{\varsigma}},{\boldsymbol{\kappa}}$ are as in [W23c, proof of Lemma 4.2.1] by [BW18, Proposition 4.6] and [W23c, Theorem 4.3.1].

### v2

@@ 1.1. Representation theory
It has been pointed out that the representation theory of $\imath$quantum groups develops very slowly ([Le19], [Wa21]). In fact, we do not still know any general theory to classify the finite-dimensional simple modules over $\imath$quantum groups of finite type, although such $\imath$quantum groups are quantum deformations of the universal enveloping algebra of a finite-dimensional complex reductive Lie algebra. Several type-dependent classification can be found in e.g. [GK91], [IK05], [M06] [IT10], [Wa20], [We20], [Wa21]. Since the $\imath$quantum groups have close resemblance to the quantum groups, it seems to be essential for better understanding of their representation theory to formulate the notion of integrable modules over $\imath$quantum groups. This is the aim of the present paper.
@@ Acknowledgments
The author thanks Paul Terwilliger for comments on the representation theory of $\imath$quantum groups, especially the $q$-Onsager algebra. He is also grateful to Stefan Kolb for fruitful discussion on a sufficient condition for a weight $\mathbf{U}^{\imath}$-module to be integrable, which stimulates him to formulate Proposition 4.3.1. This work was supported by JSPS KAKENHI Grant Number JP24K16903.
@@ 3.1. Quantum symmetric pair
Let $(I,I_{\bullet},\tau)$ be a generalized Satake diagram ([RV20, Definition 1]). From now on, we assume that the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$ is a Satake datum of type $(I,I_{\bullet},\tau)$ ([Wa23c, Definition 3.1.3]). In particular, the $I_{\bullet}$ is a subdatum of $I$ which is supposed to be of finite type, the $\tau$ is an involutive automorphism on the Cartan datum $I$, and the lattices $Y$ and $X$ are supposed to be equipped with involutive automorphisms $\tau$ which are compatible with the one on $I$. Let $w_{\bullet}\in W$ denote the longest element of the Weyl group $W_{\bullet}\subseteq W$ associated with $I_{\bullet}$. Set
@@ 3.2. Weight modules
A $\mathbf{U}^{\imath}$-module $V$ is said to be a weight module ([Wa23a, Definition 3.3.2]) if it admits a linear space decomposition
@@
as $\mathbf{U}^{\imath}$-modules. Then, the assertion can be found in [Wa23a, after Definition 3.3.5]. ∎
@@ 3.3. Integrable modules
Combining the two homomorphisms above, we see that $\mathbf{L}v^{\imath}_{\lambda,\mu}\simeq\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. By Lemma 3.3.2, the $\mathbf{U}$-module structure of $V^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}v^{\imath}_{\lambda,\mu}$. On the other hand, by [Wa23b, Proposition 3.4.3], the $\mathbf{U}$-module structure of $L^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. Hence, the assertion follows. ∎
@@ 4.1.1. Case $1:$ $\tau(k)=k=w_{\bullet}k$
Now, let us consider the $(n+1)$-dimensional irreducible representation $\mathbf{V}_{n}$ of $U_{q}(\mathfrak{sl}_{2})$, which we shall identify with the subalgebra of $\mathbf{U}$ generated by $E_{k}$, $F_{k}$, and $K_{k}^{\pm 1}$. By a similar argument to [Wa21, §3.1], one can show that the $B_{k}\in\mathbf{U}^{\imath}$ acts on $\mathbf{V}_{n}$ semisimply with eigenvalues $\kappa^{(n)},\kappa^{(n-2)},\dots,\kappa^{(-n)}$, where we set $\kappa^{(0)}:=\kappa_{k}$ and $b:=q_{k}\varsigma_{k}$. Therefore, we see that
@@ 4.2. A Presentation of the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$
$\{\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu}+\mathrm{wt}^{\imath}(b)}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}bm^{\imath}_{\lambda,\mu}\mid k\in I_{\circ},\ b\in\mathbf{B}_{\mathbf{L}}(-\infty)\}.$
Proof.
The assertion follows from Lemmas 4.2.2, 4.2.3, 4.2.4, and 4.2.5. ∎
Corollary 4.2.7.
Let $V$ be an integrable $\mathbf{U}^{\imath}$-module. Then, for each $\zeta\in X^{\imath}$ and $v\in V_{\zeta}$, there exist $(a_{j})_{j\in I_{\bullet}},(b_{j})_{j\in I_{\bullet}}\in\mathbb{Z}_{\geq 0}^{I_{\bullet}}$ and $(c_{k})_{k\in I_{\circ}}\in\mathbb{Z}_{\geq 0}^{I_{\circ}}$ such that
$E_{j}^{(a_{j}+1)}v=F_{j}^{(b_{j}+1)}v=0\ \text{ for all }j\in I_{\bullet}$
and
$\mathfrak{B}_{k,\zeta}^{(c_{k}+1)}v=0\ \text{ for all }k\in I_{\circ}.$
4.3. Some sufficient conditions
In this subsection, we give some sufficient conditions for a weight $\mathbf{U}^{\imath}$-module to be integrable.
Proposition 4.3.1.
Let $V$ be a weight $\mathbf{U}^{\imath}$-module generated by a weight vector $v$ of weight $\zeta$. Suppose that there exist $(a_{j})_{j\in I_{\bullet}},(b_{j})_{j\in I_{\bullet}}\in\mathbb{Z}_{\geq 0}^{I_{\bullet}}$ and $(c_{k})_{k\in I_{\circ}}\in\mathbb{Z}_{\geq 0}^{I_{\circ}}$ such that
$E_{j}^{(a_{j}+1)}v=F_{j}^{(b_{j}+1)}v=0\ \text{ for all }j\in I_{\bullet}$
and
$\mathfrak{B}_{k,\zeta}^{(c_{k}+1)}v=0\ \text{ for all }k\in I_{\circ}.$
Then, the $V$ is integrable.
Proof.
The assumption on $v$ implies that the $\mathbf{L}$-module $\mathbf{L}v$ is integrable. Since the $\mathbf{L}$ is a quantum group of finite type, every cyclic integrable module is finite-dimensional. Therefore, the set
$B_{V}:=\{b\in\mathbf{B}_{\mathbf{L}}(-\infty)\mid bv\neq 0\}$
is finite.
Let $\lambda,\mu\in X^{+}$ be such that $\overline{w_{\bullet}\lambda+\mu}=\zeta$. By replacing $\lambda$ and $\mu$ by $\lambda+\tau(\nu)$ and $\mu+\nu$, respectively for a suitable $\nu\in X^{+}$, we may assume that
$\langle h_{j},-w_{\bullet}\lambda\rangle\geq a_{j},$
$\langle h_{j},\mu\rangle\geq b_{j},$
$\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle\geq c_{k}$
for all $j\in I_{\bullet}$, $k\in I_{\circ}$, and $b\in B_{V}$. These imply that
$E_{j}^{(\langle h_{j},-w_{\bullet}\lambda\rangle+1)}v\in\mathbf{U}^{\imath}E_{j}^{(a_{j}+1)}v=0,$
$F_{j}^{(\langle h_{j},\mu\rangle+1)}v\in\mathbf{U}^{\imath}F_{j}^{(b_{j}+1)}v=0,$
$\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu}+\mathrm{wt}^{\imath}(b)}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}bv=b\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu}}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}v\in\mathbf{U}^{\imath}\mathfrak{B}_{k,\zeta}^{(c_{k}+1)}v=0$
for all $j\in I_{\bullet}$, $k\in I_{\circ}$, and $b\in B_{V}$. Hence, the assertion follows from Theorem 4.2.6. ∎
@@ 5. Quantized coordinate algebras
For example, the last two conditions above are satisfied if the parameters ${\boldsymbol{\varsigma}},{\boldsymbol{\kappa}}$ are as in [Wa23c, proof of Lemma 4.2.1] by [BW18, Proposition 4.6] and [Wa23c, Theorem 4.3.1].